A cylindrical frame consisting of 2 circles and 4 vertical support bars are built from a piece of wire 180cm long, cut into 6 sections: 2 circles of one length and 4 straight pieces of another length. Find the lengths of the sections that will maximize the volume of the cylinder. [Vcylinder = πr^2h]


1
Expert's answer
2021-01-26T03:27:24-0500

"\text{the length }L=180\text{ of the frame is}"

"L= 2*2\pi{r}+4h"

"\text{where }r\text{ base radius}"

"h\text{ heigth of the frame }"

"180 = 4\pi{r}+4h"

"r = \frac{45-h}{\pi}"

"V = \pi{r^2h}"

"\text{volume increases if } r>1\text{ and } h>1"

"\frac{45-h}{\pi}>1"

"h<45-\pi"

"r= \frac{45-h}{\pi};h\isin(1,45-\pi)"

Answer:"r= \frac{45-h}{\pi};h\isin(1,45-\pi)"